OpenStudy (anonymous):

Proove: a) 1+sinx=(sin(x/2)+cos(x/2))² b) sin(x+π)+sin(x-π)=-2sinx c) cosx+cos(x+120°)+cos(x+240°)=0 I tried to proove all of these but I always got something weird that I don't understand.

OpenStudy (anonymous):

Can you show how you tried?? Let us start with first one..

OpenStudy (anonymous):

You write, may be it is the case, that your approach is good, but you are doing one or two mistake in between which we will correct, so, don't worry about mistakes, just tell me how you started the first one?

OpenStudy (anonymous):

okay, just wait a minute i'll find if in my notes

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

Sure, take your time.. We are in no hurry.. :)

OpenStudy (anonymous):

1+sinx=(sin(x/2)+cos(x/2))² 1+sinx=sin²(x/2)+2sin(x/2)cos(x/2)+cos²(x/2) 1+sinx=sin(x/2)sin(x/2)+2sin(x/2)cos(x/2)+cos(x/2)cos(x/2)

OpenStudy (anonymous):

thats all I was able to do with it.

OpenStudy (anonymous):

You are just close to it..

OpenStudy (anonymous):

Basically you need to know three formulae here to do this problem, one you know, other two I will tell you.. :)

OpenStudy (anonymous):

\[1. \quad \sin^2(a) + \cos^2(a) = 1\] \[2. \quad 2\sin(a) \cdot \cos(a) = \sin(2a)\]

OpenStudy (anonymous):

you are good up to your second step..

OpenStudy (anonymous):

I mean no need to go for third step..

OpenStudy (anonymous):

Can you try to use that formulae in second step or need help?

OpenStudy (anonymous):

\[\sin^2(\frac{x}{2}) + \cos^2(\frac{x}{2}) + 2 \sin(\frac{x}{2}) \cos(\frac{x}{2})\]

OpenStudy (anonymous):

i'll try it

OpenStudy (anonymous):

Okay, you must try yourself, you will be benefited more.. :) And any doubt or any problem anywhere, let me know.. :)

OpenStudy (anonymous):

1+sinx=(sin(x/2)+cos(x/2))² 1+sinx=sin²(x/2)+2sin(x/2)cos(x/2)+cos²(x/2) 1+sinx=sin²(x/2)+cos²(x/2)+2sin(x/2)cos(x/2) sin²(x/2)+cos²(x/2)=1 2sin(x/2)cos(x/2)=sinx 1+sinx=1+sinx

OpenStudy (anonymous):

Oh wow.. :)

OpenStudy (anonymous):

But, you need to know how to represent your proof nicely.. :)

OpenStudy (anonymous):

You are writing Right Hand Side (RHS) and Left Hand Side(LHS) altogether, Take one side at the time, prove it equal to other side.. :)

OpenStudy (anonymous):

Getting me??

OpenStudy (anonymous):

well, teacher at elementary school told us that we just have to write this: left=right

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

If you want to know, then I can tell.. :)

OpenStudy (preetha):

Show him how to do it Water...

OpenStudy (anonymous):

Okay Ma'am.. :)

OpenStudy (anonymous):

i was just trying to get throught the language barrier, sorry :D

OpenStudy (anonymous):

You should take any one of the side, like you are doing it for RHS then you will do like this: